OpenStudy (anonymous):

Hi everyone! Can someone please tell me if I set up my integral correctly? I have to find the area of the inner circle. I will post my formula and answer, if someone could confirm my work that would be great! :o)

OpenStudy (anonymous):

r=2-4sin(theta) is the graph

OpenStudy (anonymous):

my integral is as follows: \[1/2\int\limits_{\pi/6}^{\pi/2}(2-4\sin \theta)^{2} d \theta \times 2\]

OpenStudy (anonymous):

my answer was \[4\pi-6\sqrt{3}\]

OpenStudy (anonymous):

When I graphed the polar curve from zero to pi/2, I basically got 1/2 of the inner circle, so I just multiplied the area formula times 2...can anyone confirm if I did this correctly? Thanks :o)

OpenStudy (anonymous):

I will leave this question open but I have to go to sleep. If anyone can confirm or deny my answer I would appreciate it! Thanks! :o)

OpenStudy (hunus):

Why is your upper limit pi/2?

OpenStudy (hunus):

That would explain why you're only getting half of the inner loop. To find out your limits you're going to have to find where the curve crosses the origin (at r=0) \[r=0=2-4 \sin (\theta)\] \[\sin(\theta)=\frac{1}{2}\] So your limits will be \[\theta = \frac{\pi}{6},\frac{5\pi}{6}\]

OpenStudy (anonymous):

hi...I will explain why I used pi/2...

OpenStudy (hunus):

But your answer is correct

OpenStudy (anonymous):

i physically graphed the curve...when theta = zero, r = 2 so your first point is (2,0)... when theta is pi/2, r = -2... the other two angles pi/4 and pi/3 simply lead you tpwards the point at angle pi/2... if you look at the entire graph, going from angle zero to pi/2 looks like it draws exactly half of the inner circle...does that make sense? hold on and i will post a picture of the graph...

OpenStudy (hunus):

Yea that makes sense. You got the correct answer lol

OpenStudy (anonymous):

Take a look at that graph...if it is wrong, can you please tell me where I messed up? :o)

OpenStudy (anonymous):

I would guess there would be a couple other ways of finding the area, but this was the first way I tried

OpenStudy (hunus):

Yea, it's correct :)

OpenStudy (anonymous):

yay! :o)

OpenStudy (anonymous):

I have a conceptual question for you if you don't mind...

OpenStudy (hunus):

Go for it :)

OpenStudy (anonymous):

if i'm not mistaken, I could have also integrated between pi/6 and pi, that way I would have had to multiply by 2... additionally, if you wanted to be really crazy, couldn't i have integrated from zero to 2pi and then subtracted the integral from pi/6 to pi right? :o)~ yikes!

OpenStudy (anonymous):

i mean NOT multiply by 2

OpenStudy (anonymous):

or heck, i think i could have even went from zero to 2pi then subtracted the (integral from pio/6 to pi/2 multiplied by 2) lol...right?

OpenStudy (hunus):

Integrating from pi/6 to pi would add a little bit too much to your area